OpenStudy (anonymous):

2log(5)=5 True or False

jhonyy9 (jhonyy9):

log base ?

OpenStudy (anonymous):

base is 2

jhonyy9 (jhonyy9):

false

OpenStudy (anonymous):

\[2\log_{2} (5)=5\]

jhonyy9 (jhonyy9):

because 5 = log(2) on exponent5

OpenStudy (anonymous):

don't i bring the 2 over to the 5, so its 5 squared.. so the answer is 25?

jhonyy9 (jhonyy9):

5 can be writing us log like log(2) on exponent5 because than this will be 5log2 because log2=1

OpenStudy (anonymous):

2*log2(5)=log2(5^2)=log2(25) which is between 4 and 5, not 5.